The makers of a certain automobile advertise that it will accelerate from 15 to 50 mi/hr in 10 s. Assume that

the acceleration is constant, Compute the distance the car travels for a period of 10 sec.


1
Expert's answer
2021-10-14T11:33:42-0400

The acceleration is


"a=\frac{\Delta v}{t}=\frac{50-15}{10}=3.5\text{ mi/(hr}·\text{s})."